Elmwood, Inc. currently sells 13,100 units of its product per year for $111 each. Variable costs total $86 per unit. Elmwood’s manager believes that if a new machine is leased for $232,525 per year, modifications can be made to the product that will increase its retail value. These modifications will increase variable costs by $14.50 per unit, but Elmwood is hoping to sell the modified units for $141 each.

a-1. Should Elmwood modify the units or sell them as is?

  • Sell as is

  • Sell with modifications

a-2. How much will the decision affect profit?
Amount =

b. What is the least Elmwood could charge for the modified units to make it worthwhile to modify them? (Round your answer to 2 decimal places.)
Minimum Selling Price =

c. The leasing company is willing to negotiate the price of the machine lease. What is the most Elmwood would be willing to pay to lease the machine if they plan to charge $141 for the modified units?
Maximum Amount =
Answer #1

Contribution margin per unit before modification = Selling price per unit- variable costs per unit

= 111-86 =$25

After modification = 141-100.5 =$40.5

Gain from modifying units = Incremental contribution margin- Incremental costs

=(40.5-25)*13100 - 232,525

=-$29,475

I.e. loss

Hence, sell as is

B.Amount =$29,475

C.maximum amount for lease = Incremental contribution margin

=$203,050 per year